# A dimer view on Fox’s trapezoidal conjecture

### v1

@@ Abstract.
Fox’s conjecture (1962) states that the sequence of absolute values of the coefficients of the Alexander polynomial of alternating links is trapezoidal. While the conjecture remains open in general, a number of special cases have been settled, some, quite recently: Fox’s conjecture was shown to hold for special alternating links by Hafner, Mészáros, and Vidinas (2023) and for certain diagrammatic Murasugi sums of special alternating links by Azarpendar, Juhász, and Kálmán (2024). Any alternating link can be written as a diagrammatic Murasugi sums of special alternating links; thus, if one could remove the certain adjective in the previous result, it would settle Fox’s conjecture. In this paper, we give an alternative proof of Azarpendar, Juhász, and Kálmán’s aforementioned result via a dimer model for the Alexander polynomial. In doing so, we not only obtain a significantly shorter proof of Azarpendar, Juhász, and Kálmán’s result than the original, but we also obtain several theorems of independent interest regarding the Alexander polynomial, which are readily visible from the dimer point of view. Moreover, we explain, backed with computational evidence, why our result is only for certain diagrammatic Murasugi sums of special alternating links as opposed to all. While the latter does not settle an open question, it strikes us as important to bring to attention, given the interest in finding a solution to Fox’s conjecture.
@@ 1. Introduction
More recently, Fox’s conjecture was settled for special alternating links by Hafner, Mészáros and Vidinas [HMV-paper1], and an alternative proof of the same result was given by Kálmán, Mészáros and Postnikov [kmp2025]. We recall that a special alternating link admits a diagram which is alternating and has only type 1 Seifert circles [^1] A Seifert circle is type 1 if it bounds a region of the link diagram, and otherwise is type 2.. Using this result as a base case, Azarpendar, Juhász and Kálmán [kalman] settled Fox’s conjecture for alternating links which admit an alternating diagram whose type 2 Seifert circles have small “length” (see Definition 4.1). We note that [kalman] use the language of “diagrammatic Murasugi sum” to describe the class of alternating links in their result.
The main object of this paper is to (1) translate Azarpendar, Juhász and Kálmán’s aforementioned result into more combinatorial language, (2) give a simpler proof of it than the original, and (3) unravel why the main idea of the proof only works when the type 2 Seifert circles are small in length (in particular, length at most 2). If this restriction on length could be removed, then Fox’s conjecture – after quite some time – would finally be settled.
To obtain our alternative proof, we use a dimer model for the Alexander polynomial similar to those used in [cohen_dissertation, TwistedDimer, perfect_matchings, cluster_alg]. Utilizing this model we give a short proof of the following theorem. See Sections 2, 3 and 4 for the relevant definitions.
@@
We also explain through computational evidence why the basic idea of the proof of Theorem 1.2 does not seem to apply when some type 2 Seifert circle has length 3 or greater. While this suggests that this approach cannot be extended to settle Fox’s conjecture, we hope that identifying its apparent pitfalls will inspire others to find a way around them.
@@
The above theorem appears as Theorem 5.15 in the text.
@@
Roadmap of the paper. In Section 2 we lay out the combinatorial dimer model that we utilize in the paper. In Section 3 we give an exposition of special alternating links, alternating links, and diagrammatic Murasugi sums. In Section 4 we translate the diagrammatic Murasugi sum into dimer language. The latter leads us to the proofs of our main results, namely Theorems 5.2,5.15, 5.16 and 5.17 in Section 5. Finally, in Section 6 we give computational evidence showing that one should not expect the proof of the main theorem to lift to arbitrary diagrammatic sums of special alternating links (i.e. to all alternating links).
@@ 2.1. The symmetrized Alexander polynomial
The symmetrized Alexander polynomial [^2] Also called the Conway normalization of the Alexander polynomial in [K06] $\widetilde{\Delta}_{L}\in\mathbb{Z}[t^{\pm 1/2}]$ is an element of the equivalence class $\Delta_{L}$ whose support is centered around zero. That is, if the highest degree term of $\widetilde{\Delta}_{L}$ is $t^{a}$, then the lowest degree term is $t^{-a}$. This determines $\widetilde{\Delta}_{L}$ up to sign; our conventions below will also fix the sign. Note that either all elements of the support of $\widetilde{\Delta}_{L}$ are in $\mathbb{Z}$ (if $\widetilde{\Delta}_{L}$ has odd degree) or all elements in the support $\widetilde{\Delta}_{L}$ are in $\mathbb{Z}+\frac{1}{2}$ (if $\widetilde{\Delta}_{L}$ has even degree). In the former case, the Alexander polynomial is saturated with respect to the lattice $\mathbb{Z}$, that is, the support of the Alexander polynomial are the integer points of an interval. In the latter case, the same is true after multiplication by $t^{1/2}$.
@@ 2.2. The face-crossing incidence graph of a link diagram
In light of Lemma 2.6, we abuse notation and, for the remainder of the paper, use $\mathcal{S}_{L,i}$ to refer to the set of perfect matchings of $\mathcal{G}_{L,i}$.
@@ 3. Special alternating links, alternating links, and the diagrammatic Murasugi sum
In this section, we define special alternating links, the diagrammatic Murasugi sum of special alternating links following Remark 2.17 of [kalman]. A key property of any alternating link is that it is a diagrammatic Murasugi sum of special alternating links [murasugi1965].
@@ 3.2. The diagrammatic Murasugi sum
In this section, we define the diagrammatic Murasugi sum of two link diagrams along a T2 circle, as in Remark 2.17 of [kalman]. A key result of Murasugi is that any alternating link is a diagrammatic Murasugi sum of special alternating links Murasugi [murasugi1965]; we use the terminology diagrammatic Murasugi sum as in [kalman].
@@
Figure 5. An illustration of diagrammatic Murasugi sum using the notation of Definition 3.2. On the left, a link diagram $L$ with a T2 circle $C$ highlighted in red and $R^{\prime\prime}$ shaded in grey. The crossings on $C$ belonging to $R^{\prime}$ are in yellow, and those belonging to $R^{\prime\prime}$ are in blue. On the right, the links $L^{\prime}$ and $L^{\prime\prime}$ so that $L=L^{\prime}*_{C}L^{\prime\prime}$. The segment $i$ illustrates 4.2, and $i^{\prime},i^{\prime\prime}$ are the unique segments of $L^{\prime},L^{\prime\prime}$ containing $i$, and are indicated in violet.
@@ 4. Notation and Setup
Figure 6. A summary of 4.3. Right: A link diagram. The T2 circle $C$ is indicated in red, and the region $R^{\prime\prime}$ bounded by $C$ is shaded grey. The groups of crossings $X_{j}$ are yellow, and the crossings $Y_{j}$ are blue. The regions $r^{\prime}_{j}$ and $r^{\prime\prime}_{j}$ are labeled. Observe $r^{\prime\prime}_{1}$ and $r^{\prime}_{3}$ are the absent regents from selecting the purple segment $i$. Left: The graph $\mathcal{G}_{L,i}$ for the link diagram on the right, with its flock edges colored pink.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$, where $C$ has length $l>0$. Choose a segment $i$ of $C$ and projection of $L$ as in 4.2. See Figure 6.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$ and we have chosen segment $i$ in $C$ as in 4.2. Let $i^{\prime}$ and $i^{\prime\prime}$ be the unique segments of $L^{\prime}$ and $L^{\prime\prime}$ which contain $i$. Delete the flock edges of $\mathcal{G}_{L,i}$. The resulting weighted graph is $\mathcal{G}_{L^{\prime},i^{\prime}}\sqcup\mathcal{G}_{L^{\prime\prime},i^{\prime\prime}}$.
@@
Next we will show that deleting certain non-flock edges of $\mathcal{G}_{L,i}$ produces the truncated face-crossing incidence graphs of two links. We will first define these links, and then state and prove the lemma. The definition of these links in the special case where $C$ has length at most $2$ originally appears in [kalman, proof of Theorem 2.20, Remark 2.24]. Our definition will use the swap move demonstrated in Figure 8.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$ has length at least 2. We use the notation from 4.3, and use $i^{\prime},i^{\prime\prime}$ to denote the unique segments of $L^{\prime},L^{\prime\prime}$ containing $i$. The link diagram $\widetilde{L^{\prime}}$ is obtained from $L^{\prime}$ by performing swap moves on $i^{\prime}$ and the segment separating $X_{j}$ from $X_{j+1}$, for $j\in[l-1]$. One of the swap moves creates a segment between the first and last crossing in $X_{1}$, which we denote by $\widetilde{i^{\prime}}$. The link diagram $\widetilde{L^{\prime\prime}}$ is obtained analogously from $L^{\prime\prime}$: first change the projection of $L^{\prime\prime}$ so that the other region adjacent to $i^{\prime\prime}$ is infinite, and then do swap moves on $i^{\prime\prime}$ and segments separating $Y_{j}$ and $Y_{j+1}$, for $j\in[l-1]$. One of these swap moves creates a segment between the first and last crossing in $Y_{l}$, which we denote by $\widetilde{i^{\prime\prime}}$.
@@ 5. Proof of main result
Let $M\in\mathcal{S}_{L,i}$. Since $\mathcal{G}_{L,i}$ is bipartite, for each $\in M_{\text{flock}}, ee$ meets precisely one black vertex and one white vertex of $\mathcal{G}_{L,i}$. Recall that, by definition, black vertices correspond to crossings of $L$ and white vertices correspond to regions. By 4.3, flock edges of $M$ satisfy the following property.
@@
This follows by Definition 5.6 and 5.8. ∎
@@
Suppose $L$ is alternating and $L=L^{\prime}*_{C}L^{\prime\prime}$. If $C$ has length $1$, then $|\widetilde{\Delta}_{L}|=|\widetilde{\Delta}_{L^{\prime}}||\widetilde{\Delta}_{L^{\prime\prime}}|$. If $C$ has length $l\geq 2$, then the right-hand sides of Theorems 5.2, 5.3 and 5.5 are non-canceling.
@@
Case 1: If $C$ has length 1, then by Lemma 5.5, $\widetilde{\Delta}_{L}=\widetilde{\Delta}_{L^{\prime}}\widetilde{\Delta}_{L^{\prime\prime}}$. By Lemma 5.7, right-hand side is non-canceling, so by Lemma 5.10,
@@
By Remarks 3.3 and 4.8, we have that $L^{\prime},L^{\prime\prime}$ are alternating and have fewer T2 circles than $L$, which are again all of length at most 2 as they are inherited from $L$. Thus by induction hypothesis $|\widetilde{\Delta}_{L^{\prime}}|$ and $|\widetilde{\Delta}_{L^{\prime\prime}}|$ are trapezoidal. By Lemma 5.14, this implies that their product $|\widetilde{\Delta}_{L^{\prime}}|\cdot|\widetilde{\Delta}_{L^{\prime\prime}}|=|\widetilde{\Delta}_{L}|$ is also trapezoidal, completing the inductive step.
@@
Unfortunately, as we will see, partitioning $\mathcal{S}^{\prime}_{L,i}$ according to flock edges used is not compatible with 6.1.
@@
as well as the complementary sums. However, none of these sums are sums over matchings of subgraphs of $\mathcal{G}_{L,i}$. So, partitioning $\mathcal{S}^{\prime}_{L,i}$ according to which flock edges are used in each matching does not help solve 6.1.
@@
The example in Figure 11 shows that the answer to 6.2 is “no.” Only 8 non-root nodes of the tree give centered trapezoidal polynomials. These nodes are

### v3

@@ Abstract.
Fox’s conjecture (1962) states that the sequence of absolute values of the coefficients of the Alexander polynomial of alternating links is trapezoidal. While the conjecture remains open in general, a number of special cases have been settled, some quite recently: Fox’s conjecture was shown to hold for special alternating links by Hafner, Mészáros, and Vidinas (2023) and for certain diagrammatic Murasugi sums of special alternating links by Azarpendar, Juhász, and Kálmán (2024). In this paper, we give an alternative proof of Azarpendar, Juhász, and Kálmán’s aforementioned beautiful result via a dimer model for the Alexander polynomial. In doing so, we not only obtain a significantly shorter proof of Azarpendar, Juhász, and Kálmán’s result than the original, but we also obtain several theorems of independent interest regarding the Alexander polynomial, which are readily visible from the dimer point of view.
@@ 1. Introduction
More recently, Fox’s conjecture was settled for special alternating links by Hafner, Mészáros and Vidinas [HMV-paper1], and an alternative proof of the same result was given by Kálmán, Mészáros and Postnikov [kmp2025]. We recall that a special alternating link admits a diagram which is alternating and has only type 1 Seifert circles [^1] A Seifert circle is type 1 if it bounds a region of the link diagram, and otherwise is type 2.. Using this result as a base case, with an beautiful insight, Azarpendar, Juhász and Kálmán [kalman] settled Fox’s conjecture for alternating links which admit an alternating diagram whose type 2 Seifert circles have small “length” (see Definition 4.1). We note that [kalman] use the language of “diagrammatic Murasugi sum” to describe the class of alternating links in their result.
The main object of this paper is to (1) translate Azarpendar, Juhász and Kálmán’s aforementioned result into the dimer language, (2) give a simpler proof of it than the original, and (3) unravel why the main idea of the proof currently only works when the type 2 Seifert circles are small in length (in particular, length at most 2). If this restriction on length could be removed, then Fox’s conjecture – after quite some time – would finally be settled.
To obtain our alternative proof, we use a dimer model for the Alexander polynomial similar to those used in [cohen_dissertation, TwistedDimer, perfect_matchings, cluster_alg]. Utilizing this model we give a short proof of the following beautiful theorem originally discovered and proved by Azarpendar, Juhász and Kálmán [kalman]. See Sections 2, 3 and 4 for the relevant definitions.
@@
We also explain through computational evidence why the basic idea of the proof of Theorem 1.2 does not seem to apply when some type 2 Seifert circle has length 3 or greater. We hope that identifying these apparent pitfalls will inspire others to find a way around them.
@@
The above theorem appears as Theorem 5.15 in the text. The length $1$ and $2$ cases of the above theorem also appear in [kalman].
@@
Roadmap of the paper. In Section 2 we lay out the combinatorial dimer model that we utilize in the paper. In Section 3 we give an exposition of special alternating links, alternating links, and diagrammatic Murasugi sums. In Section 4 we translate the diagrammatic Murasugi sum into dimer language. The latter leads us to the proofs of our main results, namely Theorems 5.2,5.15, 5.16 and 5.17 in Section 5. Finally, in Section 6 we give computations showing why our approach did not extend to all alternating links.
@@ 2.1. The symmetrized Alexander polynomial
The symmetrized Alexander polynomial [^2] Also called the Conway normalization of the Alexander polynomial in [K06] $\widetilde{\Delta}_{L}\in\mathbb{Z}[t^{\pm 1/2}]$ is an element of the equivalence class $\Delta_{L}$ whose support is centered around zero. That is, if the highest degree term of $\widetilde{\Delta}_{L}$ is $t^{a}$, then the lowest degree term is $t^{-a}$. This determines $\widetilde{\Delta}_{L}$ up to sign; our conventions below will also fix the sign. Note that either all elements of the support of $\widetilde{\Delta}_{L}$ are in $\mathbb{Z}$ (if $\widetilde{\Delta}_{L}$ has odd degree) or all elements in the support $\widetilde{\Delta}_{L}$ are in $\mathbb{Z}+\frac{1}{2}$ (if $\widetilde{\Delta}_{L}$ has even degree). Moreover, if the link $L$ is alternating, then in the former case, the Alexander polynomial $\widetilde{\Delta}_{L}$ is saturated with respect to the lattice $\mathbb{Z}$, that is, the support of the Alexander polynomial are the integer points of an interval. In the latter case, for an alternating link $L$, the same is true after multiplication by $t^{1/2}$.
@@ 2.2. The face-crossing incidence graph of a link diagram
In light of Remark 2.6, we abuse notation and, for the remainder of the paper, use $\mathcal{S}_{L,i}$ to refer to the set of perfect matchings of $\mathcal{G}_{L,i}$.
@@ 3. Special alternating links, alternating links, and the diagrammatic Murasugi sum
In this section, we define special alternating links and the diagrammatic Murasugi sum of special alternating links following [kalman, Remark 2.17]. A key property of any alternating link is that it is a diagrammatic Murasugi sum of special alternating links [murasugi1965].
@@ 3.2. The diagrammatic Murasugi sum
In this section, we define the diagrammatic Murasugi sum of two link diagrams along a T2 circle, following [kalman, Remark 2.17]. A key result of Murasugi is that any alternating link is a diagrammatic Murasugi sum of special alternating links Murasugi [murasugi1965]; we use the terminology diagrammatic Murasugi sum as in [kalman].
@@
Figure 5. An illustration of diagrammatic Murasugi sum using the notation of Definition 3.2. On the left, a link diagram $L$ with a T2 circle $C$ highlighted in red and $R^{\prime\prime}$ shaded in grey. The crossings on $C$ belonging to $R^{\prime}$ are in yellow, and those belonging to $R^{\prime\prime}$ are in blue. On the right, the links $L^{\prime}$ and $L^{\prime\prime}$ so that $L=L^{\prime}*_{C}L^{\prime\prime}$. The segment $i$ illustrates Convention 4.2, and $i^{\prime},i^{\prime\prime}$ are the unique segments of $L^{\prime},L^{\prime\prime}$ containing $i$, and are indicated in violet.
@@ 4. Notation and Setup
Figure 6. A summary of Notation 4.3. Right: A link diagram. The T2 circle $C$ is indicated in red, and the region $R^{\prime\prime}$ bounded by $C$ is shaded grey. The groups of crossings $X_{j}$ are yellow, and the crossings $Y_{j}$ are blue. The regions $r^{\prime}_{j}$ and $r^{\prime\prime}_{j}$ are labeled. Observe $r^{\prime\prime}_{1}$ and $r^{\prime}_{3}$ are the absent regents from selecting the purple segment $i$. Left: The graph $\mathcal{G}_{L,i}$ for the link diagram on the right, with its flock edges colored pink.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$, where $C$ has length $l>0$. Choose a segment $i$ of $C$ and projection of $L$ as in Convention 4.2. See Figure 6.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$ and we have chosen segment $i$ in $C$ as in Convention 4.2. Let $i^{\prime}$ and $i^{\prime\prime}$ be the unique segments of $L^{\prime}$ and $L^{\prime\prime}$ which contain $i$. Delete the flock edges of $\mathcal{G}_{L,i}$. The resulting weighted graph is $\mathcal{G}_{L^{\prime},i^{\prime}}\sqcup\mathcal{G}_{L^{\prime\prime},i^{\prime\prime}}$.
@@
Next we will show that deleting certain non-flock edges of $\mathcal{G}_{L,i}$ produces the truncated face-crossing incidence graphs of two links. We will first define these links, and then state and prove the lemma. The definition of these links in the case where $C$ has length at most $2$ originally appears in [kalman, proof of Theorem 2.20, Remark 2.24]. Our definition will use the swap move demonstrated in Figure 8.
@@
Suppose $L=L^{\prime}*_{C}L^{\prime\prime}$ has length at least 2. We use the notation from Notation 4.3, and use $i^{\prime},i^{\prime\prime}$ to denote the unique segments of $L^{\prime},L^{\prime\prime}$ containing $i$. The link diagram $\widetilde{L^{\prime}}$ is obtained from $L^{\prime}$ by performing swap moves on $i^{\prime}$ and the segment separating $X_{j}$ from $X_{j+1}$, for $j\in[l-1]$. One of the swap moves creates a segment between the first and last crossing in $X_{1}$, which we denote by $\widetilde{i^{\prime}}$. The link diagram $\widetilde{L^{\prime\prime}}$ is obtained analogously from $L^{\prime\prime}$: first change the projection of $L^{\prime\prime}$ so that the other region adjacent to $i^{\prime\prime}$ is infinite, and then do swap moves on $i^{\prime\prime}$ and segments separating $Y_{j}$ and $Y_{j+1}$, for $j\in[l-1]$. One of these swap moves creates a segment between the first and last crossing in $Y_{l}$, which we denote by $\widetilde{i^{\prime\prime}}$.
@@ 5. Proof of main result
Let $M\in\mathcal{S}_{L,i}$. Since $\mathcal{G}_{L,i}$ is bipartite, for each $\in M_{\text{flock}}, ee$ meets precisely one black vertex and one white vertex of $\mathcal{G}_{L,i}$. Recall that, by definition, black vertices correspond to crossings of $L$ and white vertices correspond to regions. By Notation 4.3, flock edges of $M$ satisfy the following property.
@@
This follows by Definition 5.6 and Notation 5.8. ∎
@@
Suppose $L$ is alternating and $L=L^{\prime}*_{C}L^{\prime\prime}$. If $C$ has length $1$, then $|\widetilde{\Delta}_{L}|=|\widetilde{\Delta}_{L^{\prime}}||\widetilde{\Delta}_{L^{\prime\prime}}|$. If $C$ has length $l\geq 2$, then the right-hand sides of Theorem 5.2, Theorem 5.3, and Lemma 5.5 are non-canceling.
@@
Case 1: If $C$ has length 1, then by Lemma 5.5, $\widetilde{\Delta}_{L}=\widetilde{\Delta}_{L^{\prime}}\widetilde{\Delta}_{L^{\prime\prime}}$. By Remark 5.7, right-hand side is non-canceling, so by Lemma 5.10,
@@
By Remark 3.3 and Lemma 4.8, we have that $L^{\prime},L^{\prime\prime}$ are alternating and have fewer T2 circles than $L$, which are again all of length at most 2 as they are inherited from $L$. Thus by induction hypothesis $|\widetilde{\Delta}_{L^{\prime}}|$ and $|\widetilde{\Delta}_{L^{\prime\prime}}|$ are trapezoidal. By Lemma 5.14, this implies that their product $|\widetilde{\Delta}_{L^{\prime}}|\cdot|\widetilde{\Delta}_{L^{\prime\prime}}|=|\widetilde{\Delta}_{L}|$ is also trapezoidal, completing the inductive step.
@@
Unfortunately, as we will see, partitioning $\mathcal{S}^{\prime}_{L,i}$ according to flock edges used is not compatible with Question 6.1.
@@
as well as the complementary sums. However, none of these sums are sums over matchings of subgraphs of $\mathcal{G}_{L,i}$. So, partitioning $\mathcal{S}^{\prime}_{L,i}$ according to which flock edges are used in each matching does not help solve Question 6.1.
[…]
@@
The example in Figure 11 shows that the answer to Question 6.2 is “no.” Only 8 non-root nodes of the tree give centered trapezoidal polynomials. These nodes are